the moment of the force F bar = i cap-j cap+k cap newton acting at the point (-1,0,1) about the point (1,1,1) is

Grade:11

1 Answers

Aksh
6 years ago
Moment of force=F vector × r vector F vector=i-j+kr vector=i+j+k-(-i+k)=2i+jMoment of force=(i-j+k)×(2i+j) =(-i+2j+3k)